OpenStudy (goformit100):

Line through the points (–2, 6) and (4, 8) is perpendicular to the line through the points (8, 12) and (x, 24). Find the value of x.

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

Get the slope of the first line... \[\huge m_1=\frac{2}{6}=\frac13\] And now \[\huge m_2=\frac{12}{x-8}=-3\] since -3 is the negative reciprocal of 1/3, making these two lines perpendicular

OpenStudy (goformit100):

Thankyou Sir